19. Find the capacitance of the Earth .the radius of the Earth is 6,400 km.By how much will the Earth`s potential chance if it receives a quantity of electricity equal to1C?


1
Expert's answer
2022-01-31T10:21:14-0500

"C=4\pi \epsilon_0R"

Put value

"C=\frac{6400\times10^3}{9\times10^9}=7.11\times10^{-4}F"

Potential changes

"Q=C\u2206V"

"∆V=\frac{1}{7.11\times10^{-4}}=1406.25V"